OpenStudy (rainbow_dashie):

Which sets of numbers are closed under addition? Choose all answers that are correct.

OpenStudy (rainbow_dashie):

A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (rainbow_dashie):

Fan, Medal, and cookies

OpenStudy (bibby):

If you add any two elements of a set and you ALWAYS get another element in the set, that set is closed under addition. For example 2+2=4 2+4=6 and 4 and 6 are part of the even number set.

OpenStudy (anonymous):

A set of elements is closed under an operation if, when you apply the operation to elements of the set, you always get another element of the set. For example, the whole numbers are closed under addition, because if you add two whole numbers, you always get another whole number - there is no way to get anything else. But the whole numbers are _not_ closed under subtraction, because you can subtract two whole numbers to get something that is not a whole number, e.g., 2 - 5 = -3 The integers are closed under multiplication (if you multiply two integers, you get another integer), but they are _not_ closed under division, since you can divide two integers to get a rational number that isn't an integer. The rationals, however, are closed under addition, subtraction, multiplication, and division. So the statement that 'the complex numbers are closed under addition' means that if you add two complex numbers together, you are guaranteed to get a complex number as the sum. Does this help?

OpenStudy (bibby):

so let's go through each set a) no 5 + 2 = 7. 7 is not in the set. b) yes c) yes d) 0+0=0

OpenStudy (rainbow_dashie):



OpenStudy (bibby):

delishus

OpenStudy (bibby):

needs moar chocolate

OpenStudy (rainbow_dashie):

Its a snikerdoodle

OpenStudy (philcua):

Make it a chocolate snickerdoodle

OpenStudy (bibby):

drizzle glaze on that

OpenStudy (philcua):

Yess that would be so good !!!!

OpenStudy (philcua):

Filled with custard @bibby

OpenStudy (bibby):

yo I'm hungry let's not talk about food right now

OpenStudy (philcua):

Ok havent had breakfast either

OpenStudy (bibby):

om nom nom

OpenStudy (philcua):

Hey bibby I dare you to go to a all you can eat buffet!!!! @bibby

OpenStudy (bibby):

lel noooooooo my mom is making food right now and im too hungover to do anything outside

OpenStudy (bibby):

omnomnomnom

OpenStudy (philcua):

Bye

OpenStudy (bibby):

you look naked